If the observed value is 8, the expected value is is 5, an the standard error is 2, then the test value used for the z test is equal to what?
If the observed value is 7, the expected value is 8, and the standard error is 2, then the test value used for the z test is equal to what?
For the conjecture "The average rent of an apartment is more than $750 per month", what is the alternative hypothesis?
Is the following a valid null hypothesis? Ho : = 6

Answer by stanbon(75887)

If the observed value is 8, the expected value is is 5, and the standard error is 2, then the test value used for the z test is equal to what?
----
z(8) = (8-5)/2 = 3/2
----------------------------
If the observed value is 7, the expected value is 8, and the standard error is 2, then the test value used for the z test is equal to what?
z(7) = (7-8)/2 = -1/2
----------------------------
For the conjecture "The average rent of an apartment is more than $750 per month", what is the alternative hypothesis?
---
Ans: The average rent is <= 750 per month.
----------------------------------------------------
Is the following a valid null hypothesis? Ho : = 6
It should say u = 6 to be valid.
Ho must be a statement about a population parameter (u or p)
======================================
Cheers,
Stan H.
============